# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v2

@@ Abstract
Understanding the geometry of the loss landscape near a minimum is key to explaining the implicit bias of gradient-based methods in non-convex optimization problems such as deep neural network training and deep matrix factorization. A central quantity to characterize this geometry is the maximum eigenvalue of the Hessian of the loss, which measures the sharpness of the landscape. Currently, its precise role has been obfuscated because no exact expressions for this sharpness measure were known in general settings. In this paper, we present the first exact expression for the maximum eigenvalue of the Hessian of the squared-error loss at any minimizer in general overparameterized deep matrix factorization (i.e., deep linear neural network training) problems, resolving an open question posed by Mulayoff & Michaeli (2020). To complement our theory, we empirically investigate an escape phenomenon observed during gradient-based training near a minimum that crucially relies on our exact expression of the sharpness.
@@ 1 Introduction
Ultimately, this understanding boils down to understanding the geometry of the loss landscape near a minimum. The maximum eigenvalue of the Hessian of the loss serves as a key measure to quantify the sharpness of the landscape near a minimum. Despite its significance, its precise role is not well-understood, particularly because closed-form expressions are generally unknown, outside a few particular cases. We summarize current state of understanding as well as the contributions of our paper in Table 1.
Table 1: Closed-form expressions for the maximum Hessian eigenvalue in the literature. $\Omega$ denotes the set of all global minimizers, $\Omega_{0}\subseteq\Omega$ denotes the set of flat global minimizers, and $\Omega_{B}\subseteq\Omega$ denotes the set of balanced global minimizers.
@@
Mulayoff & Michaeli (2020, Theorem 1) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{0}$ $\mathbb{R}^{a\times b}$
@@
In particular, in Theorem 5, we provide a closed-form expression for the maximum Hessian eigenvalue at arbitrary minima of depth-$L$ overparameterized deep matrix factorization. To the best of our knowledge, our analysis provides the first exact expression of the maximum eigenvalue for deep matrix factorization/deep linear neural network problems. In the case of deep overparameterized scalar factorization (Theorem 20) and depth-$2$ matrix factorization (Corollary 40), our closed-form expression simplifies considerably.
@@
GD always escapes from a dynamically unstable minimum, regardless of how close the initialization is to that minimum.
We explore this phenomenon through the lens of dynamical stability and, in particular, the notion of dynamically unstable minima introduced by Nar & Sastry (2018) and Wu et al. (2018). Their definition is as follows.
Definition 1 (see Wu et al. 2018, Definition 1).
Let $\bm{x}^{*}$ be a minimum of a loss function $\mathcal{L}$. Consider the linearized GD dynamics about $\bm{x}^{*}$
@@
where $\eta$ is the step size and $\nabla^{2}\mathcal{L}(\bm{x}^{*})$ is the Hessian matrix of the loss function at $\bm{x}^{*}$. Then, $\bm{x}^{*}$ is said to be dynamically unstable if, for every constant $C>0$, there exists a $t>0$ such that
$\norm{\bm{x}_{t}}^{2}_{2}>C\norm{\bm{x}_{0}}^{2}_{2}.$ (2)
We rely on the necessary and sufficient conditions established by Wu et al. (2018) for a minima $\bm{x}^{*}$ to be dynamically unstable. Their condition is that $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{x}^{*}))>2/\eta$. Thus, we see that an exact expression for the maximum eigenvalue is necessary to sharply explore the escape phenomenon.
@@ 1.1 Related Work
Balanced Minima. Ghosh et al. (2025) provided a full characterization of the Hessian spectrum at balanced minima in deep matrix factorization. Furthermore, they showed that the maximum eigenvalue of the Hessian at the flattest minima is equal to that of the balanced minima. Ding et al. (2024) showed that norm-minimal, balanced, and flat solutions coincide in depth-$2$ matrix factorization, where flatness/sharpness measured by the scaled trace of the Hessian matrix of loss function. Wang et al. (2022) showed that large step size GD training induces a balancing effect between factors in depth-$2$ matrix factorization.
Dynamical Stability. In dynamical systems theory, it is well established that asymptotic convergence to a critical point is determined solely by the local stability of that point (Strogatz, 2024). In the seminal work of Wu et al. (2018) on the dynamical stability analysis of GD training, it was shown that a global minimum is dynamically stable for GD if and only if the step size does not exceed $2/\lambda_{\max}$. Mulayoff et al. (2021) investigated this mechanism in the space of learned functions for two-layer overparameterized univariate ReLU networks in the interpolation regime. This was then extended to multivariate ReLU networks by Nacson et al. (2023). The interpolation assumption was then removed by Qiao et al. (2024); Liang et al. (2025).
Edge-of-Stability. Cohen et al. (2021) observed that neural networks trained with GD typically operate in a regime called edge of stability, in which the maximum eigenvalue of the Hessian of loss function hovers just above the value $2/\eta$, where $\eta$ is the step size, and argued that classical optimization theory fails to explain this phenomenon. Recently, Liang et al. (2025) empirically observed that explicit regularization seems to break the edge-of-stability phenomenon.
Flatness/Sharpness and Generalization. It is widely recognized in the literature that flat minima are associated with better generalization (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017). In a large-scale empirical investigation, Jiang et al. (2020) examined different measures for deep networks and found that a sharpness-based measure exhibited the strongest correlation with generalization. There is also theoretical evidence for this phenomenon in low-rank matrix recovery (Ding et al., 2024). On the other hand, Dinh et al. (2017) showed that good minima can be arbitrarily sharp in deep neural networks.
@@ 2 Notation, Preliminaries, and Problem Setup
For any $\bm{W}^{*}_{1},\bm{W}^{*}_{2},\cdots,\bm{W}^{*}_{L}]$ such that $\bm{M}=\prod_{j=1}^{L}\bm{W}^{*}_{i}$, the directional second derivative is given by
@@
is the total number of parameters in the model. $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ denotes the optimal parameters, $L\geq 2$ denotes the depth of factorization and $\bm{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ is the $i^{th}$ factor (layer). This objective is analogous to that of deep linear neural networks. To guarantee the feasibility of factorization at all points in $\mathbb{R}^{d_{L}\times d_{0}}$, we require
@@
3 Warm-Up: Deep Overparameterized Scalar Factorization
@@
Proof.
According to (12) and (13),
$=\max_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\norm{\bm{u}_{i}}_{2}\Biggr)^{2}.$ (25)
We can upper bound the right-hand side of (21) by using the triangle inequality. By applying the vectorization trick of the Kronecker product, we can rewrite (22). Then, noting the fact that for any matrix $\bm{A}\in\mathbb{R}^{m\times n}$ and vector $\bm{x}\in\mathbb{R}^{n}$, $\norm{\bm{Ax}}_{2}\leq\sigma_{\max}(\bm{A})\norm{\bm{x}}_{2}$, we can upper bound the right-hand side of (23). Note that for any matrix $\bm{A}$ and $\bm{B}$, $\sigma_{\max}(\bm{A}\otimes\bm{B})=\sigma_{\max}(\bm{A})\sigma_{\max}(\bm{B})$. Hence, we can rewrite (24).
Since $\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\norm{\bm{u}_{i}}_{2}\geq 0$ for all $\bm{u}_{i}$ such that $\sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1$, we can write the following equivalence
$\mathop{\mathrm{arg\,max}}_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\norm{\bm{u}_{i}}_{2}\Biggr)^{2}$ (26)
$=$ $\mathop{\mathrm{arg\,max}}_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\norm{\bm{u}_{i}}_{2}.$ (27)
This constrained optimization problem is a specific case of the following general constrained optimization problem.
$\min_{\bm{x}\in\mathbb{R}_{+}^{L}}-\bm{c}^{\top}\bm{x}\quad\text{s.t}\quad\norm{\bm{x}}_{2}^{2}\leq 1,$ (28)
where $\bm{c}\geq\bm{0}$. Notice that this is a convex optimization problem. Moreover, it is straightforward to verify that the optimal solution must have unit norm. To solve it, we can formulate the Lagrangian as follows:
$L(\bm{x},\mu)=-\bm{c}^{\top}\bm{x}+\mu(\bm{x}^{\top}\bm{x}-1)\quad\mu\geq 0.$ (29)
We know that optimal solution satisfies KKT conditions. Therefore,
$\bm{c}+2\mu\bm{x}^{*}=0\rightarrow\bm{x}^{*}=\frac{\bm{c}}{2\mu}\rightarrow\bm{x}^{*}=\frac{\bm{c}}{\norm{\bm{c}}_{2}}.$ (30)
If you select
$\bm{c}=\begin{bmatrix}\sigma_{\max}\Bigl(\prod_{j=2}^{L}\bm{W}^{*}_{j}\Bigr)\\ \sigma_{\max}\Bigl(\prod_{j=3}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\bm{W}_{1}^{*}\Bigr)\\ \vdots\\ \sigma_{\max}\Bigl(\prod_{j=1}^{L-1}\bm{W}^{*}_{j}\Bigr)\end{bmatrix}$ (31)
then this implies
$\phantom{{}={}}\max_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\norm{\bm{u}_{i}}_{2}\Biggr)^{2}$
$=2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{2},$ (32)
and
Then, it suffices to show that there exists a direction $\bm{U}_{1}^{*},\bm{U}_{2}^{*},\dots,\bm{U}_{L}^{*}]$ on hypersphere along which the bound in (33) is achieved.
Then,
Since the upper bound is achieved, it implies
@@ 4 Overparameterized Deep Matrix Factorization
The proof appears in Appendix B.1. Note that the deep overparameterized scalar factorization is a special case of deep matrix factorization where both $\bm{B}_{i}^{\top}\bm{B}_{i}$ and $\bm{A}_{i}\bm{A}_{i}^{\top}$ reduce to scalars. In that special case, we recover Theorem 20. Another corollary of Theorem 5 is the maximum Hessian eigenvalue for the classical (depth-$2$) matrix factorization problem. This result may be of independent interest as the expression simplifies considerably.
@@
where $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ is the optimal parameters and $\bm{L}\in\mathbb{R}^{d_{L}\times k}$, $\bm{R}\in\mathbb{R}^{d_{0}\times k}$. To ensure the feasibility of the factorization every point in $\mathbb{R}^{d_{L}\times d_{0}}$, we choose $k\geq\min\{d_{0},d_{L}\}$. We define the set of minimizers as follows:

### v3

@@ Abstract
Understanding the geometry of the loss landscape near a minimum is key to explaining the implicit bias of gradient-based methods in non-convex optimization problems such as deep neural network training and deep matrix factorization. A central quantity to characterize this geometry is the maximum eigenvalue of the Hessian of the loss, which measures the sharpness of the landscape. Currently, its precise role has been obfuscated because no exact expressions for this sharpness measure were known in general settings. In this paper, we present the first exact expression for the maximum eigenvalue of the Hessian of the squared-error loss at any minimizer in general overparameterized deep matrix factorization (i.e., deep linear neural network training) problems, resolving an open question posed by Mulayoff & Michaeli (2020). This expression uncovers a fundamental property of the loss landscape of depth-2 matrix factorization problems: a minimum is flat if and only if it is spectral-norm balanced, which implies that flat minima are not necessarily Frobenius-norm balanced. Furthermore, to complement our theory, we empirically investigate an escape phenomenon observed during gradient-based training near a minimum that crucially relies on our exact expression of the sharpness.
@@ 1 Introduction
Ultimately, this understanding boils down to understanding the geometry of the loss landscape near a minimum. The maximum eigenvalue of the Hessian of the loss serves as a key measure to quantify the sharpness of the landscape near a minimum. Despite its significance, its precise role is not well-understood, particularly because closed-form expressions are generally unknown, outside a few particular cases. We summarize the current state of understanding as well as the contributions of our paper in Table 1.
Table 1: Closed-form expressions for the maximum Hessian eigenvalue in the literature. $\Omega$ denotes the set of all global minimizers, $\Omega_{F}\subseteq\Omega$ denotes the set of flat global minimizers, and $\Omega_{B}\subseteq\Omega$ denotes the set of balanced global minimizers.
@@
Mulayoff & Michaeli (2020, Theorem 1) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{F}$ $\mathbb{R}^{a\times b}$
@@
In particular, in Theorem 5, we provide a closed-form expression for the maximum Hessian eigenvalue at arbitrary minima of depth-$L$ overparameterized deep matrix factorization. We also highlight that overparameterized deep matrix factorization and the deep linear neural network setting of Mulayoff & Michaeli (2020) are equivalent for investigating the sharpness measure $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ when the data covariance matrix is identity, i.e., $\bm{\Sigma}_{x}=\bm{I}$ (see Mulayoff & Michaeli, 2020, Equation 21). To the best of our knowledge, our analysis provides the first exact expression of the maximum eigenvalue for deep matrix factorization/deep linear neural network problems. In the case of deep overparameterized scalar factorization (Theorem 21) and depth-$2$ matrix factorization (Corollary 25), our closed-form expression simplifies considerably.
@@
GD always escapes from a dynamically unstable minimum.
We explore this phenomenon through the lens of dynamical stability introduced by Wu et al. (2018).
Consider a twice continuously differentiable loss function $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$, and let $\bm{x}^{*}$ be a minimizer of $\mathcal{L}$. For $\delta>0$, denote the open ball $B_{\delta}(\bm{x}^{*}):=\left\{\bm{x}\in\mathbb{R}^{N}:\norm{\bm{x}-\bm{x}^{*}}_{2}<\delta\right\}$. There exists a $\delta>0$ such that, for all $\bm{x}\in B_{\delta}(\bm{x}^{*})$,
$\mathcal{L}(\bm{x})=\mathcal{L}(\bm{x}^{*})+(\bm{x}-\bm{x}^{*})^{\top}\nabla\mathcal{L}(\bm{x}^{*})+\frac{1}{2}(\bm{x}-\bm{x}^{*})^{\top}\nabla^{2}\mathcal{L}(\bm{x}^{*})(\bm{x}-\bm{x}^{*})+o(\norm{\bm{x}-\bm{x}^{*}}_{2}^{2}).$ (1)
Definition 1.
Let $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$ be a twice continuously differentiable loss function, and let $\bm{x}^{*}$ be a minimizer of $\mathcal{L}$. The corresponding linearized GD update rule for $\mathcal{L}$ is given by
@@
where $\eta>0$ is the step size. Given $\delta>0$ such that (1) holds, if there exists an initial point $\bm{x}_{0}\in B_{\delta}(\bm{x}^{*})$ such that the residuals $\bm{\epsilon}_{t}=\bm{x}_{t}-\bm{x}^{*}$ diverge, i.e.,
$\lim_{t\rightarrow\infty}\norm{\bm{\epsilon}_{t}}_{2}=\infty,$ (3)
then $\bm{x}^{*}$ is said to be dynamically unstable.
A necessary and sufficient condition for a minimum to be dynamically unstable is that $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{x}^{*}))>2/\eta$ (cf. Wu et al. (2018); Mulayoff et al. (2021); Chemnitz & Engel (2025)). Thus, we see that an exact expression for the maximum eigenvalue is necessary to explore the escape phenomenon empirically.
1.1 Contributions
In this paper, we present the first exact expression for the maximum eigenvalue of the Hessian of the squared-error loss at any minimizer in general overparameterized deep matrix factorization problems. Our results lead to the following remarkable observations:
  • A minimizer of deep overparameterized scalar factorization loss is flat if and only if the product of spectral norms of left and right intermediate factors is constant across layers (Corollary 7).
  • Flat minima are spectral-norm balanced in depth-2 matrix factorization (Corollary 33). This implies that flat minima are not necessarily balanced, contrary to claims made in several works (Ding et al., 2024; Ghosh et al., 2025). We further discuss this in Section 6.
  • A minimizer of deep matrix factorization loss is flat if the product of spectral norms of left and right intermediate factors is constant across layers (Corollary 10).
  • A recent work (Anonymous, 2025) provides empirical evidence that, in two-layer linear networks, the largest eigenvalue of the Hessian of the loss function is correlated with the spectral norm of the weight matrices. Furthermore, for deep linear networks, they observe that this largest Hessian eigenvalue correlates with the spectral norm of the products of the weight matrices. Our results in Theorem 21, Theorem 5, and Corollary 25 provide the first theoretical explanation for their empirical observations.
@@ 1.2 Related Work
Balanced Minima. Ghosh et al. (2025) provided a full characterization of the Hessian spectrum at balanced minima in deep matrix factorization. Furthermore, they showed that the maximum eigenvalue of the Hessian at the flattest minima is equal to that of the balanced minima. Ding et al. (2024) showed that norm-minimal, balanced, and flat solutions coincide in depth-$2$ matrix factorization, where sharpness measured by the scaled trace of the Hessian matrix of the loss function. We further discuss Ghosh et al. (2025) and Ding et al. (2024)’s results in Section 6. Finally, Wang et al. (2022) showed that large step size GD training induces a balancing effect between factors in depth-$2$ matrix factorization.
Dynamical Stability. In dynamical systems theory, it is well established that asymptotic convergence to a critical point is determined solely by the local stability of that point (Strogatz, 2024). In the seminal work of Wu et al. (2018) on the dynamical stability analysis of GD training, it was shown that a global minimum is dynamically stable for GD if and only if the step size does not exceed $2/\lambda_{\max}$, where $\lambda_{\max}$ denotes the maximum eigenvalue of the Hessian of the loss at the minimum. Mulayoff et al. (2021) investigated this mechanism in the space of learned functions for two-layer overparameterized univariate ReLU networks in the interpolation regime. This was then extended to multivariate ReLU networks by Nacson et al. (2023). The interpolation assumption was then removed by Qiao et al. (2024); Liang et al. (2025).
Edge-of-Stability. Cohen et al. (2021) observed that neural networks trained with GD typically operate in a regime called edge of stability, in which the maximum eigenvalue of the Hessian of the loss function hovers just above the value $2/\eta$, where $\eta$ is the step size, and argued that classical optimization theory fails to explain this phenomenon. Recently, Liang et al. (2025) empirically observed that explicit regularization seems to break the edge-of-stability phenomenon.
Sharpness and Generalization. It is widely recognized in the literature that flat minima are associated with better generalization (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017). In a large-scale empirical investigation, Jiang et al. (2020) examined different measures for deep networks and found that a sharpness-based measure exhibited the strongest correlation with generalization. There is also theoretical evidence for this phenomenon in low-rank matrix recovery (Ding et al., 2024). On the other hand, Dinh et al. (2017) showed that good minima can be arbitrarily sharp in deep neural networks.
@@ 2 Notation, Preliminaries, and Problem Setup
For any $\bm{W}^{*}_{1},\bm{W}^{*}_{2},\cdots,\bm{W}^{*}_{L}]$ such that $\bm{M}=\prod_{j=1}^{L}\bm{W}^{*}_{j}$, the directional second derivative is given by
@@
is the total number of parameters in the model. $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ denotes the matrix contains the parameters subject to factorization, $L\geq 2$ denotes the depth of factorization and $\bm{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ is the $i^{th}$ factor (layer). This objective is analogous to that of deep linear neural networks. To guarantee the feasibility of factorization at all points in $\mathbb{R}^{d_{L}\times d_{0}}$, we require
@@
Local Minima. Laurent & Brecht (2018) showed that all local minima of deep linear networks with convex and differentiable loss are global if the layers satisfy (17), i.e., hidden layers are at least as wide as input and output layers. In particular, they proved a more general theorem concerning real-valued functions that take as input a product of matrices. Thus, a corollary of Laurent & Brecht (2018, Theorem 1) is that overparameterized deep matrix factorization does not have spurious minima, i.e., all local minima are global.
3 Deep Overparameterized Scalar Factorization
@@
The proof appears in Appendix B.
@@ 4 Overparameterized Deep Matrix Factorization
The proof appears in Appendix C.1. Note that the deep overparameterized scalar factorization is a special case of deep matrix factorization where both $\bm{B}_{i}^{\top}\bm{B}_{i}$ and $\bm{A}_{i}\bm{A}_{i}^{\top}$ reduce to scalars. In that special case, we recover Theorem 21. Another corollary of Theorem 5 is the maximum Hessian eigenvalue for the classical (depth-$2$) matrix factorization problem. This result may be of independent interest as the expression simplifies considerably.
@@
where $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ is the target matrix and $\bm{L}\in\mathbb{R}^{d_{L}\times k}$, $\bm{R}\in\mathbb{R}^{d_{0}\times k}$. To ensure the feasibility of the factorization every point in $\mathbb{R}^{d_{L}\times d_{0}}$, we choose $k\geq\min\{d_{0},d_{L}\}$. We define the set of minimizers as follows: